Alexander runs his own canteen. This morning, he sold 13 "chef's plates" and 32 "quick and delicious" snacks. He made $143.25 on these sales. In the afternoon, he sold 7 "chef's plates" and 23 "quick and delicious" snacks, and collected $93. Use this information to find the cost of each meal.
Answer by Cromlix(4381)

Hi there,
Make 'chef's plates' = x
and 'Quick and Delicious snacks' = y
13x + 32y = 143.25.........(1)
7x + 23y = 93..............(2)
Multiply Eq(1) by 7 and Eq(2) by 13
91x + 224y = 1002.75.......(1)
91x + 299y = 1209..........(20
Subtract Eq(1) from Eq(2)
......75y = 206.25
......y = 2.75
Substitute y = 2.75 into
Eq(2)
7x + 23y = 93..............(2)
7x + 23(2.75) = 93
7x + 63.25 = 93
7x = 93 - 63.25
7x = 29.75
x = 4.25
Chef's plates = $4.25
Quick and Easy Snacks = $2.75
Hope this helps :-)